OpenStudy (anonymous):

Solving quadratic equations, Solve 2x2 - 8x = -7.

OpenStudy (causticsyndicalist):



OpenStudy (causticsyndicalist):

or +, either way works.

OpenStudy (anonymous):

x = 4 ± √ 2/2

OpenStudy (causticsyndicalist):

yeah, that ^

OpenStudy (anonymous):

I cant seem to be able to read that @heavenfun13

OpenStudy (anonymous):

reload the page

OpenStudy (anonymous):

\[-2 \pm \sqrt{2}\] \[-2 \pm 2\sqrt{2}\] \[\frac{ 2\pm \sqrt{2} }{ 2 }\] \[2\pm \frac{ \sqrt{2} }{ 2 }\] This is what I have for answers to pick from @heavenfun13 @CausticSyndicalist

OpenStudy (causticsyndicalist):

It's the second one I believe.

OpenStudy (causticsyndicalist):

or last one.

OpenStudy (anonymous):

3rd